# Scalable and robust regression models for continuous proportional data

### v1

@@ Abstract
Beta regression is used routinely for continuous proportional data, but it often encounters practical issues such as a lack of robustness of regression parameter estimates to misspecification of the beta distribution. We develop an improved class of generalized linear models starting with the continuous binomial (cobin) distribution and further extending to dispersion mixtures of cobin distributions (micobin). The proposed cobin regression and micobin regression models have attractive robustness, computation, and flexibility properties. A key innovation is the Kolmogorov-Gamma data augmentation scheme, which facilitates Gibbs sampling for Bayesian computation, including in hierarchical cases involving nested, longitudinal, or spatial data. We demonstrate robustness, ability to handle responses exactly at the boundary (0 or 1), and computational efficiency relative to beta regression in simulation experiments and through analysis of the benthic macroinvertebrate multimetric index of US lakes using lake watershed covariates.
@@ 1 Introduction
Regression analysis of proportional data, corresponding to bounded continuous data in the unit interval $[0,1]$, is a common focus in many fields. Such data arise in diverse contexts, ranging from the analysis of rates or indices in econometrics (Papke and Wooldridge,, 1996) to the percentage of tissue area in medical imaging (Peplonska et al.,, 2012). We are particularly motivated by ecological applications; Warton and Hui, (2011) estimates $\sim 14\%$ of ecology articles involve data based on proportions that are not derived from counts, with examples including measurements of percent coverage (Korhonen et al.,, 2024) and ecological indices (Lindholm et al.,, 2021).
@@
However, beta regression has several limitations. First, beta distributions are not in the natural exponential family (Morris,, 1982) and thus do not strictly belong to the GLM class. This implies that the mean and dispersion parameters are not orthogonal to each other (Ferrari and Cribari-Neto,, 2004), and the well-established properties of GLMs may not hold. Beta regression also faces computational difficulties when considering extensions to accommodate complex dependence structures. This limitation is particularly relevant in Bayesian frameworks, which support flexible hierarchical extensions (Bolker et al.,, 2009). Finally, the presence of exact 0s or 1s prevents the direct application of beta regression models. This issue is often bypassed by manipulating the data to be within the open interval (0,1) (Smithson and Verkuilen,, 2006), but the results are often highly sensitive to such preprocessing (Kosmidis and Zeileis,, 2024).
@@
The motivation of this article is to introduce a proper GLM approach to continuous proportional response data, without the need for data preprocessing and facilitating computation, including in complex settings involving random effects. With these goals in mind, we propose continuous binomial (cobin) regression, a name inspired by models based on the continuous Bernoulli distribution (Loaiza-Ganem and Cunningham,, 2019; Quijano Xacur,, 2019). The cobin distribution is an exponential dispersion model (Jørgensen,, 1987) with orthogonal mean and dispersion parameters, implying that the corresponding GLM has appealing properties, including consistency of maximum likelihood estimates (MLE) under model misspecification. For posterior sampling using the canonical link function, we propose a novel data augmentation strategy based on Kolmogorov-Gamma random variables, leading to a conditionally normal likelihood, similar to Pólya-Gamma augmentation for logistic models (Polson et al.,, 2013). We develop an efficient sampler for Kolmogorov-Gamma random variables. The proposed Gibbs sampler can easily accommodate extensions to Gaussian latent variables and random effects, including in spatial contexts. Furthermore, we show uniform ergodicity of the proposed MCMC algorithms, providing theoretical guarantees on efficient inference procedure.
@@
We also propose micobin regression, an extension of cobin regression based on dispersion mixtures of continuous binomial (micobin) distributions. This improves flexibility and robustness by localizing the dispersion parameter (Wang and Blei,, 2018), and can handle exact 0s or 1s without modifying the data. Micobin regression addresses non-structural boundary values (Blasco-Moreno et al.,, 2019) and is distinct from models that explicitly handle boundary values by assigning positive probability mass to 0s or 1s, such as zero/one inflated or censored models (Ospina and Ferrari,, 2012; Kubinec,, 2023; Kosmidis and Zeileis,, 2024). We describe how micobin regression can be further extended to allow the dispersion parameter to systematically vary according to covariates. Through simulations, we show that micobin regression achieves better predictive performance than beta regression in misspecified settings.
@@
The paper is structured as follows. In Section 2, we introduce cobin and micobin regressions, study their properties, and present hierarchical extensions. In Section 3, we establish the Kolmogorov-Gamma integral identity in Theorem 1, which could be of independent interest for other models, and describe our Kolmogorov-Gamma augmentation strategy for posterior computation. In Section 4, we develop a highly efficient algorithm to sample Kolmogorov-Gamma random variables. In Section 5 and 6, we perform simulation studies to support our claims and present an application example. All proofs of theorems are in Appendix S.1 except for Theorem 1.
@@ 2.1 Continuous binomial distribution
Our goal is to find a family of distributions supported on the unit interval that serve as the random component of a GLM having appealing properties. Unlike binary regression where the conditional mean completely determines the distribution, it is desirable to have an additional dispersion parameter instead of just the natural parameter. We introduce the continuous binomial (cobin) distribution, which is in the exponential dispersion family and contains the uniform distribution as a special case. We first define the density of the cobin distribution. To avoid overloading notation, in what follows, we assume that $(e^{x}-1)/x$, $\sinh(x)/x$, and similar expressions are well defined in the limit at $x=0$.
@@
When $\theta=0$ and $\lambda=1$, the cobin distribution reduces to $\mathrm{Unif}(0,1)$. In Appendix S.3, we provide a formal derivation of the cobin as an exponential dispersion model, including a proof that $\lambda$ must be an integer, along with comparable derivations for the normal, gamma, and inverse Gaussian families. The cobin distribution has appeared in the literature in various contexts, dating back to Bates, (1955) in the study of time intervals between accidents. The special case of $\lambda=1$, with different parameterizations based on $\varphi=1/(1+e^{-\theta})$, is called the continuous Bernoulli distribution (Loaiza-Ganem and Cunningham,, 2019), since the density is proportional to $\varphi^{y}(1-\varphi)^{1-y}$. The name “continuous binomial” comes from the fact that the cobin distribution arises from the convolution of i.i.d. continuous Bernoulli random variables, $Y_{1},\dots,Y_{\lambda}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}\mathrm{cobin}(\theta,1)$ then $\lambda^{-1}\sum_{l=1}^{\lambda}Y_{l}\sim\mathrm{cobin}(\theta,\lambda^{-1})$, which follows from the convolution property of the exponential dispersion model (Jørgensen,, 1987).
@@
which illustrates that $E(Y)$ is only controlled by the natural parameter $\theta$ and $\operatorname*{var}(Y)$ is proportional to the dispersion parameter $\lambda^{-1}$. See Figure 2 for examples and a comparison with the beta distribution having the same mean and variance.
An important distinction from the beta distribution is that the cobin distribution does not allow bimodal densities having spikes at both zero and one. This implies that the range of possible variances under the cobin distribution is smaller than that for that beta, and the maximum variance $1/12$ is achieved at $\mathrm{Unif}(0,1)$. This restriction may be an advantage in many applications in which the true density of the response given covariates is unlikely to be U-shaped, but U-shaped fitted distributions can arise in beta regression for certain covariate values due to lack of fit and sparse data.
@@ 2.2 Cobin regression
With a link function $g:(0,1)\to\mathbb{R}$ that is strictly monotone and differentiable, the GLM with continuous binomial response can be expressed as:
$Y_{i}\mid\theta_{i},\lambda\stackrel{{\scriptstyle\mathrm{ind}}}{{\sim}}\mathrm{cobin}(\theta_{i},\lambda^{-1}),\quad\theta_{i}=(B^{\prime})^{-1}\{g^{-1}(\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta})\},\quad i=1,\dots,n,$ (2)
which implies $g\{E(y_{i}\mid x_{i})\}=\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta}$. We refer to the model (2) as cobin regression. Denoting the linear predictor $\eta_{i}=\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta}$ and the conditional mean $\mu_{i}=g^{-1}(\eta_{i})$, the likelihood equations for $\bm{\beta}$ are
$\frac{\partial}{\partial\beta_{j}}\sum_{i=1}^{n}\log p_{\mathrm{cobin}}(y_{i};\theta_{i},\lambda^{-1})=\lambda\sum_{i=1}^{n}\frac{(y_{i}-\mu_{i})x_{ij}}{B^{\prime\prime}(\theta_{i})}\frac{\partial\mu_{i}}{\partial\eta_{i}}=0,\quad j=1,\dots,p$ (3)
where the solution $\hat{\bm{\beta}}$ does not depend on the dispersion parameter $\lambda^{-1}$. Standard procedures, such as Newton–Raphson or iteratively reweighted least squares, can be used to find the MLE as well as maximum a posteriori (MAP) estimates.
Unlike beta regression, where two shape parameters corresponding to sufficient statistics $T(y)=\{\log(y),\log(1-y)\}$ jointly control the mean, cobin regression is a natural exponential family with a natural parameter $\theta$ corresponding to sufficient statistics $T(y)=y$. Therefore, unlike beta regression, cobin regression is a proper GLM and hence inherits many attractive properties, such as a concave log-likelihood function under the canonical link. We highlight the robustness property of the solution of (3) against model misspecification.
@@
Under mild regularity conditions (Gourieroux et al.,, 1984), $\hat{\bm{\beta}}$ of the cobin regression likelihood equations (3) is consistent if $E(y_{i}\mid x_{i})=g^{-1}(\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta})$ is correctly specified.
Since the cobin distribution with fixed $\lambda$ belongs to the natural exponential family, Proposition 1 is a direct consequence of viewing (3) as quasi-maximum likelihood estimating equations associated with a natural exponential family (Gourieroux et al.,, 1984); also see §4.2.6 and §8.3 of Agresti, (2015) for details. Thus, under the correct mean structure, the cobin regression produces an asymptotically valid point estimate even if $Y$ does not follow the cobin distribution.
@@
Another important difference between beta and cobin is the form of the score function, the derivative of the log-likelihood with respect to $\bm{\beta}$. In contrast to the cobin score function (3) which is bounded in terms of $y$, the beta score function depends on $\log\{y/(1-y)\}$ and is thus unbounded for $y\in(0,1)$ (Ferrari and Cribari-Neto,, 2004). Thus, $y$ near its boundary values can lead to an arbitrarily large change in the beta regression log-likelihood landscape. Given the fact that boundedness of the score function often serves as a necessary condition for robustness properties (e.g., Prop. 2 of Cantoni and Ronchetti, (2001)), cobin regression has greater robustness to observations near the boundaries than beta regression.
@@
The canonical link function of cobin regression, which we call the cobit link, satisfies $g_{\mathrm{cobit}}^{-1}(\eta)=B^{\prime}(\eta)=e^{\eta}/(e^{\eta}-1)-\eta^{-1}$. Although $g_{\mathrm{cobit}}(\mu)$ does not have a closed form expression, it can be easily inverted numerically in practice. Compared to logistic function $1/(1+e^{-\eta})$, $g_{\mathrm{cobit}}^{-1}(\eta)$ slowly approaches 0 and 1 as $|\eta|\to\infty$, asymptotically at the same speed as the Cauchy distribution c.d.f. See Figure S.5 in Appendix S.3 for details. Similarly to heavy-tailed link functions in binary response regression, this behavior is appealing in allowing the mean parameter to approach zero or one more slowly for extreme values of the predictors, reducing sensitivity to certain types of outliers.
@@
Under the canonical link function, the cobin regression model can be expressed as $Y_{i}\mid\bm{x}_{i}\stackrel{{\scriptstyle\mathrm{ind}}}{{\sim}}\mathrm{cobin}(\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta},\lambda^{-1})$ for $i=1,\dots,n$. In the following sections, we focus on cobin regression with canonical link functions, which greatly simplifies model fitting. Regardless of the choice of the link function, relationships between the conditional mean and covariates can be reported in terms of marginal effects (Williams,, 2012).
2.3 Mixture of cobin regression and varying dispersion model

### v2

@@ Abstract
Beta regression is used routinely for continuous proportional data, but it often encounters practical issues such as a lack of robustness to misspecification of the beta distribution and sensitivity to outliers. We develop an improved class of generalized linear models starting with the continuous binomial (cobin) distribution and further extending to dispersion mixtures of cobin distributions (micobin). The proposed cobin regression and micobin regression models have attractive robustness, computation, and flexibility properties. A key innovation is the Kolmogorov-Gamma data augmentation scheme, which facilitates Gibbs sampling for Bayesian computation, including in hierarchical cases involving nested, longitudinal, or spatial data. We demonstrate robustness, ability to handle responses exactly at the boundary (0 or 1), and computational efficiency relative to beta regression in simulation experiments and through analysis of the benthic macroinvertebrate multimetric index of US lakes using lake watershed covariates.
@@ 1 Introduction
Regression analysis of proportional data, corresponding to bounded continuous data in the unit interval $[0,1]$, is a common focus in many fields. Such data arise in diverse contexts, including rates or indices in economics (Brazier et al., , 2002), percentage of tissue area in medical imaging (Peplonska et al., , 2012), and rating scales for Alzheimer’s disease (Rosen et al., , 1984). We are particularly motivated by ecological applications; Warton and Hui, (2011) estimates $\sim 14\%$ of ecology articles involve data based on proportions that are not derived from counts, with examples including measurements of percent coverage (Korhonen et al., , 2024) and ecological indices (Lindholm et al., , 2021).
@@
However, beta regression has several limitations. First, the beta distribution is neither in the natural exponential family (Morris, , 1982), nor part of the broader family of exponential dispersion models (Jørgensen, , 1987), and thus beta regression does not strictly belong to the GLM class. This implies that the mean and dispersion parameters are not orthogonal to each other (Ferrari and Cribari-Neto, , 2004), and the well-established properties of GLMs may not hold. In particular, beta regression is known to be highly sensitive to outliers (Bayes et al., , 2012). Beta regression also faces computational difficulties when considering extensions to accommodate complex dependence structures. This limitation is particularly relevant in Bayesian frameworks, which support flexible hierarchical extensions (Bolker et al., , 2009). Finally, the presence of exact 0s or 1s prevents the direct application of beta regression models. This issue is often bypassed by manipulating the data to be within the open interval (0,1) (Smithson and Verkuilen, , 2006), but the results are often highly sensitive to such preprocessing (Kosmidis and Zeileis, , 2025).
@@
The motivation of this article is to introduce a proper GLM approach to continuous proportional response data, without the need for data preprocessing and facilitating computation, including in complex settings involving random effects. With these goals in mind, we propose continuous binomial (cobin) regression, a name inspired by models based on the continuous Bernoulli distribution (Loaiza-Ganem and Cunningham, , 2019; Quijano Xacur, , 2019). The cobin distribution is an exponential dispersion model (Jørgensen, , 1987) with orthogonal mean and dispersion parameters, implying that the corresponding GLM has appealing properties, including consistency of maximum likelihood estimates (MLE) under distributional misspecification. Importantly, the exponential dispersion structure also leads to bounded score functions, which provides improved robustness against boundary-proximate outliers. For posterior sampling using the canonical link function, we propose a novel data augmentation strategy based on Kolmogorov-Gamma random variables, leading to a conditionally normal likelihood, similar to Pólya-Gamma augmentation for logistic models (Polson et al., , 2013). We develop an efficient sampler for Kolmogorov-Gamma random variables. The proposed Gibbs sampler can easily accommodate extensions to Gaussian latent variables and random effects, including in spatial contexts. Furthermore, we show uniform ergodicity of the proposed MCMC algorithms, providing theoretical guarantees for efficient inference procedures.
@@
We also propose micobin regression, an extension of cobin regression based on dispersion mixtures of continuous binomial (micobin) distributions. This improves flexibility and robustness by localizing the dispersion parameter (Wang and Blei, , 2018), and can handle exact 0s or 1s without modifying the data. Micobin regression handles non-structural boundary values, such as those introduced by measurement precision limits or rounding during preprocessing, and is distinct from models that explicitly model boundary values by assigning positive probability mass to 0s or 1s, such as zero/one inflated or censored models (Ospina and Ferrari, , 2012; Kubinec, , 2023; Kosmidis and Zeileis, , 2025). We describe how micobin regression can be further extended to allow the dispersion parameter to systematically vary according to covariates. Through simulations, we show that micobin regression achieves better predictive performance under distributional misspecification and greater robustness to outliers compared to beta regression and regression models based on mixtures of beta distributions (Hahn, , 2008; Bayes et al., , 2012).
@@
The paper is structured as follows. In Section 2, we introduce cobin and micobin regressions, study their properties, and present hierarchical extensions. In Section 3, we establish the Kolmogorov-Gamma integral identity in Theorem 1, which could be of independent interest for other models, and describe our Kolmogorov-Gamma augmentation strategy for posterior computation. In Section 4 and 5, we perform simulation studies to support our claims and present an application example. All proofs of theorems are in Supplementary Section S.2.
@@ 2.1 Continuous binomial distribution
Our goal is to find a family of distributions supported on the unit interval that serve as the random component of a GLM having appealing properties. Unlike binary regression where the conditional mean completely determines the distribution, it is desirable to have an additional dispersion parameter instead of just the natural parameter. We introduce the continuous binomial (cobin) distribution, which is in the exponential dispersion family and contains the uniform distribution as a special case. We first define the density of the cobin distribution. To avoid overloading notation, we adopt the convention that $(e^{0}-1)/0\coloneqq\lim_{x\to 0}(e^{x}-1)/x=1$ and $\sinh(0)/0\coloneqq\lim_{x\to 0}\sinh(x)/x=1$, and we similarly fill in removable singularities in all functions that are otherwise infinitely differentiable on all of $\mathbb{R}$.
@@
When $\theta=0$ and $\lambda=1$, the cobin distribution reduces to $\mathrm{Unif}(0,1)$. In Supplementary Section S.3, we provide a formal derivation of the cobin as an exponential dispersion model, including a proof that $\lambda$ must be an integer, along with comparable derivations for the normal, gamma, and inverse Gaussian families. To the best of our knowledge, the earliest reference to the continuous binomial distribution appears in Bates, (1955) in the context of modeling time intervals between accidents, but has received little attention in the literature. The special case of $\lambda=1$, with reparametrization via $\varphi=1/(1+e^{-\theta})$, has been rediscovered under the name continuous Bernoulli distribution (Loaiza-Ganem and Cunningham, , 2019), since the density is proportional to $\varphi^{y}(1-\varphi)^{1-y}$. The name “continuous binomial” comes from the fact that the cobin distribution arises from the convolution of i.i.d. continuous Bernoulli random variables, $Y_{1},\dots,Y_{\lambda}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}\mathrm{cobin}(\theta,1)$ then $\lambda^{-1}\sum_{l=1}^{\lambda}Y_{l}\sim\mathrm{cobin}(\theta,\lambda^{-1})$, which follows from the convolution property of the exponential dispersion model (Jørgensen, , 1987).
@@
which illustrates that $E(Y)$ is only controlled by the natural parameter $\theta$ and $\operatorname*{var}(Y)$ is proportional to the dispersion parameter $\lambda^{-1}$; if $\theta=0$ we have $E(Y)=1/2$ and $\operatorname*{var}(Y)=1/(12\lambda)$. See Figure 2 for examples and a comparison with the beta distribution having the same mean and variance.
@@ 2.2 Cobin regression
With $\bm{x}_{i}=(x_{ij})_{j=1}^{p}\in\mathbb{R}^{p}$ and a link function $g:(0,1)\to\mathbb{R}$ that is strictly monotone and differentiable, the GLM with continuous binomial response can be expressed as:
$Y_{i}\mid\theta_{i},\lambda,\bm{x}_{i}\stackrel{{\scriptstyle\mathrm{ind}}}{{\sim}}\mathrm{cobin}(\theta_{i},\lambda^{-1}),\quad\theta_{i}=(B^{\prime})^{-1}\{g^{-1}(\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta})\},\quad i=1,\dots,n,$ (2)
which implies $g\{E(y_{i}\mid\bm{x}_{i})\}=\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta}$ for $\bm{\beta}\in\mathbb{R}^{p}$. We refer to the model (2) as cobin regression. Denoting the linear predictor $\eta_{i}=\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta}$ and the conditional mean $\mu_{i}=g^{-1}(\eta_{i})$, the likelihood equations for $\bm{\beta}$ are
$\frac{\partial}{\partial\beta_{j}}\sum_{i=1}^{n}\log p_{\mathrm{cobin}}(y_{i};\theta_{i},\lambda^{-1})=\sum_{i=1}^{n}\frac{(y_{i}-\mu_{i})x_{ij}}{\lambda^{-1}B^{\prime\prime}(\theta_{i})}\frac{\partial\mu_{i}}{\partial\eta_{i}}=0,\quad j=1,\dots,p$ (3)
where the solution $\hat{\bm{\beta}}$ does not depend on the dispersion parameter $\lambda^{-1}$. Standard procedures, such as Newton–Raphson or iteratively reweighted least squares, can be used to find the MLE $\hat{\bm{\beta}}$.
Unlike the beta distribution, where inference on the mean parameter is based on the sufficient statistics $\{\sum_{i=1}^{n}\log(y_{i}),\sum_{i=1}^{n}\log(1-y_{i})\}$ rather than $\sum_{i=1}^{n}y_{i}$, the cobin distribution belongs to an exponential dispersion model, and cobin regression is therefore a proper GLM (McCullagh and Nelder, , 1989, §2.2.2). Thus, it inherits many attractive properties of GLMs, such as a concave log-likelihood function under the canonical link. We highlight the robustness property of its MLE against model misspecification, which implies that cobin regression under the correct mean structure produces an asymptotically valid point estimate even if $Y$ does not follow the cobin distribution.
@@
Under mild regularity conditions (Gourieroux et al., , 1984), $\hat{\bm{\beta}}$ of the cobin likelihood equations (3) is consistent if $E(y_{i}\mid\bm{x}_{i})=g^{-1}(\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta})$ is correctly specified.
As shown in Gourieroux et al., (1984), obtaining a consistent estimator of $\bm{\beta}$ under potential distributional misspecification requires that $\hat{\bm{\beta}}$ is the solution to estimating equations associated with a natural (linear) exponential family. This condition is both necessary and sufficient; see also §4.2.6 and §8.3 of Agresti, (2015). Thus, Proposition 1 is a direct consequence of the cobin distribution belonging to an exponential dispersion model, where the dispersion parameter disappears in the estimating equations, satisfying the sufficient condition for consistency. Furthermore, the necessity of the condition implies that the MLE of beta regression is generally not consistent, unless the full distribution is correctly specified.
Remark 1.
Quasi-likelihood approaches, which specify only the conditional mean and variance of $Y$ without requiring a full distributional form, also yield consistency results analogous to Proposition 1. A well-known example is the fractional logit or fractional probit models based on quasi-MLE (Papke and Wooldridge, , 1996), whose quasi-score equation replaces the $\lambda^{-1}B^{\prime\prime}(\theta_{i})$ term in equation (3) with $\sigma^{2}\mu_{i}(1-\mu_{i})$ for some $\sigma^{2}>0$. However, due to a lack of proper probabilistic formulation, quasi-likelihood approaches face challenges for uncertainty quantification and hierarchical extensions analogous to mixed-effects model settings, and therefore will not be pursued further here.
Another important difference between the beta and cobin distributions is the form of the score function, the derivative of the log-likelihood with respect to $\bm{\beta}$. In contrast to the cobin regression score function (3) which is bounded in terms of $y$, the beta regression score function for a single observation $y_{i}$ and the $j$th component of $\bm{\beta}$ is
$\frac{\partial}{\partial\beta_{j}}\log p_{\mathrm{beta}}\left(y_{i};g^{-1}(\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta}),\phi\right)=\phi\left\{\log\left(\frac{y_{i}}{1-y_{i}}\right)-\frac{\Gamma^{\prime}(\mu_{i}\phi)}{\Gamma(\mu_{i}\phi)}+\frac{\Gamma^{\prime}(\phi-\mu_{i}\phi)}{\Gamma(\phi-\mu_{i}\phi)}\right\}x_{ij}\frac{\partial\mu_{i}}{\partial\eta_{i}},$ (4)
@@
where $p_{\mathrm{beta}}\left(y_{i};g^{-1}(\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta}),\phi\right)$ corresponds to the beta density evaluated at $y_{i}$ with mean-precision parametrization (Ferrari and Cribari-Neto, , 2004), and $\Gamma$ is a gamma function. From (4), the beta regression score function depends directly on $\log\{y/(1-y)\}$, which is unbounded for $y\in(0,1)$. As a result, $y$ near its boundary values can lead to an arbitrarily large change in the beta regression log-likelihood landscape. Given the fact that boundedness of the score function often serves as a necessary condition for robustness properties (e.g., Prop. 2 of Cantoni and Ronchetti, (2001)), cobin regression has greater robustness to observations near the boundaries than beta regression.
@@
The canonical link function of cobin regression, which we call the cobit link, has inverse $g_{\mathrm{cobit}}^{-1}(\eta)=B^{\prime}(\eta)=e^{\eta}/(e^{\eta}-1)-\eta^{-1}$. Although $g_{\mathrm{cobit}}(\mu)$ does not have a closed-form expression, it can be easily calculated in practice by solving $g_{\mathrm{cobit}}^{-1}(x)=\mu$ numerically. Compared to logistic function $1/(1+e^{-\eta})$, $g_{\mathrm{cobit}}^{-1}(\eta)$ slowly approaches 0 and 1 as $|\eta|\to\infty$, asymptotically at the same speed as the Cauchy c.d.f. See Supplementary Section S.3 for details. Similarly to heavy-tailed link functions in binary response regression, this behavior is appealing in allowing the mean parameter to approach zero or one more slowly for extreme values of the predictors, reducing sensitivity to certain types of outliers.
@@
Under the canonical link function, the cobin regression model can be expressed as $Y_{i}\mid\bm{x}_{i}\stackrel{{\scriptstyle\mathrm{ind}}}{{\sim}}\mathrm{cobin}(\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta},\lambda^{-1})$ for $i=1,\dots,n$. In the following sections, we focus on cobin regression with canonical link functions, which greatly simplifies model fitting. Regardless of the choice of the link function, relationships between the conditional mean and covariates can be reported in terms of average slopes, also known as average marginal effects (Williams, , 2012). For example, the average slope of a continuous covariate $X_{j}$ under the link function $g$ is computed as $\frac{1}{n}\sum_{i=1}^{n}\frac{\partial E(y\mid\bm{x})}{\partial x_{j}}\big|_{\bm{x}=\bm{x}_{i}}=\frac{1}{n}\sum_{i=1}^{n}(g^{-1})^{\prime}(\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta})\beta_{j}.$ For a categorical covariate, the partial derivative is replaced by a unit difference.
2.3 Micobin regression and varying dispersion model